### v1

# Gravitational waves from dark domain walls

Abstract
For most of cosmic history, the evolution of our Universe has been governed by the physics of a ‘dark sector’1, consisting of dark matter and dark energy, whose properties are only understood in a schematic way. The influence of these constituents is mediated exclusively by the force of gravity, meaning that insight into their nature must be gleaned from gravitational phenomena. The advent of gravitational-wave astronomy2, 3, 4 has revolutionised the field of black hole astrophysics, and opens a new window of discovery for cosmological sources. Relevant examples include topological defects, such as domain walls or cosmic strings5, 6, which are remnants of a phase transition. Here we present the first simulations of cosmic structure formation in which the dynamics of the dark sector introduces domain walls as a source of stochastic gravitational waves in the late Universe7, 8. We study in detail how the spectrum of gravitational waves is affected by the properties of the model, and extrapolate the results to scales relevant to the recent evidence for a stochastic gravitational wave background4. Our relativistic implementation9, 10, 11 of the field dynamics paves the way for optimal use of the next generation of gravitational experiments12, 13, 14, 15 to unravel the dark sector.
keywords:
N-body simulation – dark energy – domain walls – gravitational waves – gravity
The detection of the nanohertz stochastic gravitational wave background by the NANOGrav pulsar timing array4 opens a new window for studying gravitational systems at cosmological scales. The observed signal consists of correlated timing residuals of 68 millisecond pulsars that have now been measured over a 15-year time period. The amplitude of the timing residuals scales as $h_{c}^{2}f^{-3}$, where $f$ is the frequency and $h_{c}$ is the characteristic strain of the stochastic gravitational wave background16, i.e. the characteristic fractional change in the length of a hypothetical detector arm. An astrophysical background from supermassive black hole inspirals is predicted17 to yield $h_{c}\sim f^{-2/3}$ and hence a spectrum of timing residuals that tilts as $f^{-13/3}$. While supermassive black holes remain a viable explanation of the observed signal, a tilt between $f^{-2.6}$ and $f^{-3.8}$ is preferred by the data, which is in mild tension with the prediction. This has prompted a large number of investigations into alternative explanations, exploring possible phenomena in the dark sector or in the early Universe18, 19.
One hypothetical source of gravitational waves are topological defects that are generically produced at phase transitions. Depending on the nature of the spontaneously broken symmetry, the defects can take the form of cosmic strings or domain walls20, 21. In particular, if the vacuum manifold splits into two (or more) isolated points, domain walls separate the regions that have settled into different vacua during the phase transition. The case where the phase transition happens in the early Universe at a high energy scale has been studied thoroughly in the recent literature22, 23, 24, 25, 19, owing to the dimensional argument that frequencies in the nanohertz range today correspond to the size of the causal horizon (the scale that sets the typical size of different vacuum domains) when the Universe had a temperature of several trillion degrees Kelvin.
In this work instead, we consider domain walls appearing in the late-time Universe due to phase transitions in the dark energy field. This scenario has not yet been studied in the context of stochastic gravitational waves. We consider the (a)symmetron model26, a generalisation of the symmetron7, 8, a theory of gravity with an extra scalar degree of freedom. Such extra degrees of freedom are a general prediction of physics beyond the Standard Model. The (a)symmetron scalar field couples to the matter density and therefore mediates a fifth force that affects cosmological structure formation and is tightly constrained by local gravity experiments. When the local matter density falls below some critical threshold $\rho_{*}$, it triggers a $\mathbb{Z}_{2}$ phase transition, and the field settles into either of two minima. A crucial difference from the early Universe is that the late-time Universe has very large fluctuations in the local density, so that the phase transition happens in different places at different times. The resulting domains are separated by domain walls that are moulded by the sheets and filaments in the cosmic large-scale structure. This way, a large number of domain walls can be formed in a Hubble volume. Both their energy density and Compton scale are free parameters, but the former is often chosen well below the critical density such that the expansion rate of the Universe is not affected significantly. However, they can still produce a considerable stochastic gravitational wave signal if their time evolution is violent enough. In the case where the potential is not symmetric, the asymmetry tends to destabilise the domains that occupy the higher minimum. Spontaneous reconfiguration of domains then leads to the collision and annihilation of domain walls, adding a further source of gravitational waves that is qualitatively different from stable domain walls.
The asymmetron model depends on four parameters10: the Compton wavelength $L_{C}$, indicating the correlation length of the scalar field in vacuum; the cosmological redshift of the phase transition $z_{*}$, indicating the time when the background matter density is equal to the critical density $\rho_{*}$ of the phase transition; the average strength of the fifth force relative to gravity, $\bar{\beta}=(\beta_{+}+\beta_{-})/2$; and the respective difference $\Delta\beta=\beta_{+}-\beta_{-}$ in the two minima. The asymmetron reduces to the symmetron for the choice $\Delta\beta=0$, in which case the scalar potential is symmetric. We run two suites of simulations varying parameters about their respective values in a reference model, using Roman numerals \@slowromancapi@-\@slowromancapv@ for simulations in a cubic volume of $(742\,\mathrm{Mpc})^{3}$, and numerals i-v for the ones using a smaller volume of $(148\,\mathrm{Mpc})^{3}$. For quick reference, we use a subscript to indicate the parameter that was varied with respect to the reference model.
[…]
The left panel of figure 1 shows the intensity of gravitational waves emanating from collapsing domain walls for the model \@slowromancapiii@${}_{\Delta\beta}$ that has an asymmetry of $10\%$ in the scalar field potential. In contrast, the right panel shows the gravitational wave pattern for a standard $\Lambda$-cold-dark-matter ($\Lambda$CDM) cosmological model without the scalar field. In this case, the main source of long-wavelength gravitational waves is the evolving large-scale structure.
Figure 2: Power spectra of the time derivative of the tensor perturbation, $\dot{h}_{ij}$, plotted as dimensionless energy density per logarithmic frequency interval. The conversion from Fourier wavenumber $k$ to frequency $f$ assumes the standard dispersion relation $k=2\pi f/c$. Results from the simulations in the larger volume, $(742\,\mathrm{Mpc})^{3}$, are shown in the left panel and the ones for the smaller volume, $(148\,\mathrm{Mpc})^{3}$, are shown in the right panel. The dotted line labelled “LSS” indicates the gravitational wave contribution from the evolution of large-scale structure as predicted from second-order perturbation theory. The short blue line labelled “SMBH”, inserted as a visual guide only, shows the expected slope for an astrophysical signal produced by supermassive black hole inspirals. This signal would however appear at very different frequencies and amplitudes.
Figure 2 shows the present-day ($z=0$) power spectra of gravitational waves, presented in units of dimensionless energy density per logarithmic frequency interval. There is a clear generation of power at the scale corresponding to the Compton wavelength of the theory with a power-law tail extending to lower frequencies. At extremely low frequency, below about $10^{-16}\,\mathrm{Hz}$, cosmic large-scale structure (LSS) becomes the dominant source of gravitational waves in most cases. The corresponding amplitude can be predicted from second-order perturbation theory and is indicated as dotted line in the figure, showing good agreement with our simulations in the regime where perturbation theory is valid. The somewhat larger amplitude we find for $\Lambda$CDM is due to higher-order nonlinearities and, at high wavenumber, discretisation noise in our numerical scheme. We checked that our results for the scalar field models are numerically converged.
The following qualitative trends can be seen in our suite of simulations: a smaller Compton wavelength $L_{C}$ reduces the amplitude and moves the peak power to a correspondingly smaller scale; a larger average fifth force $\bar{\beta}$ increases the amplitude; and a later time of symmetry breaking (lower $z_{*}$) reduces the amplitude and has otherwise a small effect. Finally, adding asymmetry to the potential by choosing $\Delta\beta>0$, we find more unstable domains, significantly increasing the contribution of collapsing domain walls to the gravitational wave signal. We find that this component, which consists of a noise of gravitational wave bursts, has a nearly flat spectrum at low frequencies.
The NANOGrav collaboration has presented confidence intervals for the spectral tilt by fitting a power law $\sim Af^{-\gamma}$ to the timing residuals4. We measure the spectral tilt $\gamma$ from our simulations, using the fact that the timing residuals scale as $\propto\Omega_{\mathrm{gw}}f^{-5}$. Our results are shown in figure 3. We note that all simulations display a plateauing spectral tilt that persists until the Compton scale. Furthermore, there is no strong indication for a dependence on $\bar{\beta}$ and $z_{*}$. Stable domain walls seem to generate slightly bluer spectra (lower $\gamma$) compared to the NANOGrav measurement, whereas collapsing and annihilating domain walls (models with $\Delta\beta>0$) push the tilt all the way to the red side of the NANOGrav posterior.
Figure 3: The spectral index $\gamma$ inferred from the power spectra shown in figure 2. The blue bands show the 68% and 95% confidence intervals reported for the NANOGrav observations that were however taken at much higher frequencies. The dotted horizontal line labelled SMBH indicates the expected spectral index for supermassive black hole inspirals, which would also appear at much higher frequencies.
To connect our results with the NANOGrav measurements, we need to extrapolate eight orders of magnitude towards smaller scales and about ten orders of magnitude in energy density $\Omega_{\mathrm{gw}}$. Such a huge extrapolation in parameter space comes with several caveats that we address partially, and they should be investigated further in the future. To get some analytic insight, it is useful to consider that the surface tension $\sigma$ of the domain walls scales as $\bar{\beta}^{2}(1+z_{*})^{6}L_{C}^{3}$ in vacuum. This sets the strength of the gravitational wave source and we expect that $\Omega_{\mathrm{gw}}\propto\sigma^{2}$. Our simulations indeed confirm this scaling with $\bar{\beta}$, while we find a slightly stronger-than-expected dependence $\Omega_{\mathrm{gw}}\sim(1+z_{*})^{14}$. The dependence on $L_{C}$ seems much more complicated than the naive consideration of $\sigma$ suggests. This may be due to the fact that domain walls tend to be aligned with sheets and filaments of dark matter, affecting the effective potential and hence the surface tension in a non-trivial way. We point out that the Compton scale $L_{C}$ sets the typical length scale over which the field can respond to local variations in the matter density.
Given that $L_{C}$ would need to be reduced by some eight orders of magnitude from our reference simulations to reach NANOGrav frequencies, it is difficult to predict the amplitude robustly. The trends seen in our simulations suggest that the amplitude would need to be boosted significantly to reach NANOGrav levels, which can be achieved by increasing $\bar{\beta}$ or $z_{*}$, or both, presumably by many orders of magnitude. This would yield a model with a very strong fifth force that is however also strongly screened. Allowing for asymmetry in the potential ($\Delta\beta>0$) is an interesting scenario to explore, as we have seen that the gravitational wave amplitude from collapsing domain walls can easily dominate the signal.
In conclusion, we have shown that a late-time cosmological source may be a viable candidate for providing the observed spectral tilt of the NANOGrav experiment, although matching the signal amplitude remains an open question that requires a significant extrapolation of our results. Nevertheless, this work presents for the first time a full simulation of an ultra-light scalar field in a late-time cosmological $N$-body code, including a consistent treatment of gravity and sourcing of gravitational waves. Our code provides a new framework and opens the door to studies of the late-time production of gravitational waves from cosmological sources, such as topological defects, late-time phase transitions, and their interaction with structure formation. This tool is set to play an important role at the interface between cosmological surveys and gravitational-wave astronomy, both of which are needed to probe the dark sector and advance our understanding of the nature of gravity.
Methods
We write the space-time line element of the expanding Friedmann universe with linearised gravitational waves as
$ds^{2}=a^{2}(\tau)\left[-c^{2}d\tau^{2}+\left(\delta_{ij}+h_{ij}\right)dx^{i}dx^{j}\right]\,,$ (1)
where $a$ denotes the scale factor, $c$ is the speed of light, $d\tau=a^{-1}dt$ is a conformal time element, $x^{i}$ are comoving Cartesian coordinates on the spatial hypersurface, and $h_{ij}$ is the transverse and traceless tensor perturbation containing the gravitational waves. A sum is implied over repeated indices. Our simulations also track scalar and vector perturbations of the metric11, but these are relevant mainly for matter dynamics and shall be omitted in the following discussion for brevity. In Fourier space, the tensor perturbations obey a damped oscillator equation that is driven by the spin-2 projection of the stress-energy tensor27,
$\ddot{\tilde{h}}_{ij}+2\mathcal{H}\dot{\tilde{h}}_{ij}+c^{2}k^{2}\tilde{h}_{ij}=\frac{16\pi G}{c^{2}}\left(P_{i}^{l}P_{j}^{m}-\frac{1}{2}P_{ij}P^{lm}\right)\tilde{T}_{lm}\,,$ (2)
where a dot denotes a derivative with respect to $\tau$, $\mathcal{H}=d\ln a/d\tau$ is the conformal Hubble rate, $G$ is Newton’s gravitational constant, $k$ is the comoving Fourier wavenumber, and $P_{ij}=\delta_{ij}-k^{-2}k_{i}k_{j}$ is a transverse projector. A tilde indicates spatial Fourier transform, and $\tilde{T}_{lm}$ is the spatial part of the stress-energy tensor in Fourier space.
In our numerical simulations, we compute the stress-energy tensor of the scalar field, and hence of the domain walls, non-perturbatively by solving the full equations of motion discretised on a three-dimensional grid. The stress-energy is taken to Fourier space, using fast Fourier transforms, to source the wave equation (2) that we integrate separately for each wave vector using a temporal leapfrog integrator. We evolve only the Fourier components for which our time stepping guarantees a stable integration, which requires $k\lesssim(c\Delta\tau)^{-1}$ for an integration step $\Delta\tau$. At any point in time, we can estimate the power spectrum $P_{\dot{h}}(k)$ of $\dot{h}_{ij}$,
$\sum_{i,j}\left\langle\dot{\tilde{h}}_{ij}(\boldsymbol{k},\tau)\dot{\tilde{h}}_{ij}(\boldsymbol{k}^{\prime},\tau)\right\rangle=\delta^{(3)}_{\mathrm{D}}(\boldsymbol{k}+\boldsymbol{k}^{\prime})P_{\dot{h}}(k)\,,$ (3)
where $\delta^{(3)}_{\mathrm{D}}$ is the three-dimensional Dirac delta distribution. Assuming that the gravitational waves propagate freely at the time when we perform the measurement, this spatial two-point correlation is directly related to the temporal one via the retarded time. In this case one finds that the spectral density in the frequency domain, $S_{h}(f)$, is simply given by $2\pi c^{3}S_{h}(f)=P_{\dot{h}}(k=2\pi f/c)$.
It is common practice to express the spectrum in terms of the energy density of gravitational waves28,
[…]
One may furthermore define the dimensionless density per logarithmic frequency interval,
$\Omega_{\mathrm{gw}}(f)=\rho_{\mathrm{crit}}^{-1}\frac{d\rho_{\mathrm{gw}}}{d\ln f}=\frac{4\pi^{2}}{3H^{2}_{0}}f^{3}S_{h}(f)=\frac{2\pi^{2}}{3H^{2}_{0}}f^{2}h_{c}^{2}(f)\,,$ (5)
where $H_{0}$ is the present-day Hubble parameter and the critical density of the Universe is given by $\rho_{\mathrm{crit}}=3c^{2}H_{0}^{2}/(8\pi G)$.
When viewed on scales larger than the thickness of the domain wall, its stress-energy is simply proportional to the surface tension
$\sigma\simeq\int_{\phi_{-}}^{\phi_{+}}\sqrt{2V_{\rm{eff}}(\phi)-2V_{\rm{eff}}(\phi_{\pm})}\,,$ (6)
where one integrates between the two asymptotic vacuum values ϕ±subscriptitalic-$\phi_{\pm}$ on the two sides of the wall. At vanishing matter density we have $\sigma=12\sqrt{2}\,\Omega_{\mathrm{m}}^{2}H_{0}^{4}\bar{\beta}^{2}(1+z_{*})^{6}L_{C}^{3}$.
We have run [^1] The code is publicly available and can be downloaded at https://github.com/oyvach/AsGRD. three high-resolution simulations ($1280^{3}$ grid points, $742$ Mpc box size), labelled \@slowromancapi@-\@slowromancapiii@, and two lower-resolution simulations ($512^{3}$ grid points, $742$ Mpc box size), \@slowromancapiv@ and \@slowromancapv@. An additional five simulations roughly match the higher resolution but use a smaller volume ($256^{3}$ grid points, $148$ Mpc box size), and are labelled i-v. Among the larger volume simulations, the reference model \@slowromancapi@ has parameters ($L_{C},z_{*},\bar{\beta},\Delta\beta)=(1480\,\mathrm{kpc},0.1,8,0$). Models \@slowromancapii@-\@slowromancapv@ vary one parameter at a time, using $L_{C}=1111\,\mathrm{kpc}$, $\Delta\beta/\bar{\beta}=10\%$, $z_{*}=0$ and $\bar{\beta}=12$, respectively. The varied parameter is indicated as subscript in our figure labels for easy identification. In the smaller volume simulations, our reference model i has ($L_{C},z_{*},\bar{\beta},\Delta\beta)=(743\,\mathrm{kpc},-0.2,90,0$), and models ii-v use $L_{C}=801\,\mathrm{kpc}$, $\Delta\beta/\bar{\beta}=30\%$, $z_{*}=-0.13$ and $\bar{\beta}=120$, respectively.
Acknowledgements
Ø.C. thanks Constantinos Skordis and the Central European Institute of Cosmology (CEICO) for hosting him during the work on the project. J.A. acknowledges support from the Swiss National Science Foundation. We thank the Research Council of Norway for their support. The simulations were performed on resources provided by UNINETT Sigma2 – the National Infrastructure for High Performance Computing and Data Storage in Norway. This work is supported by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID s1051.
Author contributions
Ø.C. implemented the model in the code, ran the simulations and did most of the analysis. J.A. proposed and contributed to the code to obtain the gravitational wave signal, and played a leading role in the theoretical interpretation of the results. All authors discussed the results and contributed to the writing.
Competing interests
The authors declare no competing interests.
Materials & Correspondence
Correspondence and requests for materials should be directed to Øyvind Christiansen (oyvind.christiansen@astro.uio.no).